OpenStudy (anonymous):

find all solutions to cos(2x)=cosx

OpenStudy (mertsj):

Replace cos(2x) with 2cos^2x-1

OpenStudy (anonymous):

Hint: Use the double-angle formula for cos(2x) = 2 cos^2 (x) - 1

OpenStudy (anonymous):

Set all equal to zero, and factor that quadratic....

OpenStudy (mertsj):

Wow!! Wish I would have thought of that.

OpenStudy (anonymous):

Thaats why you come to this room!

OpenStudy (noelgreco):

Factor it like you would factor a quadratic, even though it's not a quadratic.

OpenStudy (anonymous):

It IS a quadratic!

OpenStudy (anonymous):

2 cos^2(x) - 1 = cos x 2 cos^2(x) - cos x - 1 = 0 (2 cos x + 1)(cos x -1) = 0 cos x = - 1/2 and cos x = 1

OpenStudy (anonymous):

I assume you wanted to solve for cos x, and not angles for x, as you haven't given interval for angles, etc.

OpenStudy (noelgreco):

Sorry, a quadratic is a second degree expression of in one variable. In this case cos x is not a variable, but a function. Again, it can be factored like a quadratic, but it is not one.

OpenStudy (anonymous):

It is a quadratic equation.

OpenStudy (anonymous):

Makes no difference if it is x, y, cos x, or whatever you want to call it.

OpenStudy (noelgreco):

Yes, it does make a difference. A quadratic is a second-degree POLYNOMIAL expression. In teaching precalculus, I made your mistake for years until corrected by a visiting professor. 2 cos^2(x) - cos x - 1 is a "quadratic type" expression, but not a quadratic. Is it a fine point? Yes. Is mathematics chock-full of fine points? Yes, again.